# Michael Callen Miguel Fajardo-Steinhäuser Michael G. Findley

### v2

@@ 1 Cash-based transfers in humanitarian contexts
In all, we identified 23 relevant cash-based transfer programs summarized in Table A10. Peer-reviewed academic studies account for 16 of the 23, five are academic working papers, and two are evaluation reports. Column 1 reports the country in which the cash-based intervention took place and the countries are ordered by the current level of fragility of the country in which they occur, moving from the least fragile (top) to most fragile (bottom) (OECD, 2022). Ecuador and Philippines appear at the top because neither country is listed in the OECD’s list of fragile countries. The next two, Lebanon and Sri Lanka, are not on the core OECD list either, but are flagged as showing key early-warning signs. Sometimes the target beneficiaries (Column 2) were refugees from another country, but located in a host country, and sometimes the intended beneficiaries were local to a country. Regardless, target beneficiaries were high need or otherwise vulnerable. Column 3 reports the type of cash-based assistance, specifically identifying cash (C) and vouchers (V), and the most relevant alternative, food assistance (F). Some of these programs include other/combined treatments, which we note in the extended discussions. Column 4 reports the delivery mechanism, whether some form of digital, physical, or a combination.
@@ 2 Work related to outcomes: Usage, food security, well-being, diversion
To contextualize our experimental findings on usage, food security and mental well-being, Table A10 reports whether the 23 reviewed studies empirically examined usage at any level, denoting examination with a check mark (✓), the same (or similar) outcome categories that we did, including whether they found positive, null, or negative effects. Few studies carried out systematic tests for diversion, so we only include a check mark (✓) if the study made any observation about levels of diversion in their context. To supplement the summary table, Table LABEL:LiteratureFindings1 provides details about specific study context and findings.
@@
Because tech-illiterate populations are largely poor, rural, elderly, and female (Peachey, 2020), it is important to assess whether digital transfers will reach those most vulnerable and be used effectively. If they cannot, then digital cash-based approaches may hurt intended beneficiaries relative to what may have occurred with physical cash or in-kind food distribution. In India’s PMGKY program intended to provide COVID relief, for example, the government transferred benefits to the accounts of over 200 million beneficiaries, but being female, illiterate, and living in a household without a smartphone resulted in extremely low use of digital payments, which ranged from 1% to 3.9% (Gentilini, 2022). Of the 10 studies in our review with a digital component, five relied at least in part on mobile platforms, meaning that the beneficiary needed to use the phone for receipt/withdrawal of funds. The remaining five relied primarily on ATM cards, which are less demanding technologically, and also confer fewer advantages such as decentralization and transparency. Of these five, only one reported on usage, noting that most of the time the funds were withdrawn immediately and, where not, they were used up eventually (Lyall, Zhou, and Imai, 2020). The remaining four did not report usage explicitly, but descriptions were suggestive of substantial usage, though we acknowledge that this assessment may be overly generous. Our own results in one of the most challenging environments, and with an arguably more vulnerable demographic than most other studies, are encouraging, though we emphasize the importance of user-centered design in our registration and training procedures. Near 100% usage by a largely illiterate and vulnerable population in a context of extreme oppression and insecurity suggests genuine potential for technology to “flatten access” to otherwise marginalized populations (Grossman, Humphreys, and Sacramone-Lutz, 2014). (See Table A10, Column 5).
@@
Most of the studies (21 of 23) considered a food security outcome in some form, although there is considerable variation in the indicators each examines. Of the 21, 16 find that cash-based transfers improved food security outcomes, five found no relationship, and none of the studies found any negative effects. Of the 21, 20 examined at least one indicator of food quantity (e.g., food expenditures or calorie consumption). Of the 21, 18 examined at least one indicator of food quality (e.g., dietary diversity). And 16 of 21 considered at least one indicator of negative coping strategies (e.g., meal skipping). These patterns add support to a growing body of evidence documenting that humanitarian cash-based programs have positive impacts on a variety of food security outcomes. Existing studies most frequently considered basic needs outcomes. In a recent systematic review, eight of 20 studies considered food security (primarily food consumption) and four of 20 studies examined dietary diversity, most of which demonstrated positive impacts (Jeong and Trako, 2022). Moreover, five studies examined reductions in negative coping strategies (e.g., meal skipping) and three found positive impacts. (See Table A10, Column 6).
@@
In contrast to food security outcomes, A10 shows that limited attention has been given to the possible mental benefits (or costs) of cash-based programs. Only seven of the 23 studies considered some mental well-being outcomes and six found positive effects. An evaluation of a voucher program in DRC found a positive impact (+0.32sd) on an overall mental well-being index at the end of six weeks, which attenuated to some extent, but remained remarkably durable at the end of one year (+0.18sd and statistically significant) (Quattrochi et al., 2022). They also measured life satisfaction, where they found a positive impact at six weeks (18% higher in treatment; control: 3.29 on 10-pt scale), but the result disappeared by the one-year mark. Perhaps the strongest impact of cash on mental well-being, with documented short-, medium-, and long-term effects of cash relative to control, occurred in Niger (Bossuroy et al., 2022). Importantly, like the other treatment arms, cash was accompanied by coaching, savings, training, and market facilitation, making the overall bundle difficult to compare to other cash-only programs. The dearth of studies examining mental well-being mirrors that of prominent systematic reviews (Jeong and Trako, 2022). They only identified two (of 20) studies reporting subjective well-being results: the DRC voucher study that we just discussed (Quattrochi et al., 2022) and a study of Syrian refugees in Jordan (that did not meet our inclusion criteria) that found isolated improvement on a single measure of self-esteem, but notably no effect on a measure of satisfaction, the closest to our life satisfaction indicator. Similarly, in another systematic review, there were no rigorously evaluated studies that addressed mental well-being though some mixed-methods evaluations have found isolated positive impacts (Attah et al., 2016; Doocy and Tappis, 2016; Hagen-Zanker et al., 2017). (See Table A10, Column 7).
@@
The extent of diversion in these studies is essentially unknown as four of the 23 studies addressed diversion, but only three attempted to measure its extent. Studies measured the amount of the cash and vouchers actually received and found little leakage (Aker, 2017), how much of the grant recipients had to give to other household and community members and participants who reported less than 1% of the grant (Blattman et al., 2016), and the occurrence of robbery where participants reported no occurrence (Lehmann and Masterson, 2020). Another study mentioned that the implementer feared diversion, and therefore hired security to guard cash during transport, which comprised the largest share of their costs, but did not report measuring diversion (Aker et al., 2016). Many policy reports have downplayed concerns about diversion, suggesting that diversion risks are likely overstated and that cash is no less prone to diversion than other types of aid (Sossouvi, 2013; Gordon, 2015; Overseas Development Institute, 2015; Ecker, Maystadt, and Guo, 2019). As far as we are aware, no systematic tests of diversion in a cash-based context have been undertaken. In our study, we experimentally tested for diversion, both using direct and indirect questions, and found no evidence of its presence. We even interviewed key stakeholders, such as merchants, to understand alternative points where diversion could have occurred, and found little indication. This finding is especially informative given how extensively the scholarly and policy literatures have identified diversion as a core concern for in-kind assistance. Indeed, anecdotal evidence of diversion has been extensively documented across the globe (Perlez, 1992; Bryer and Cairns, 1997; De Waal, 1997; Anderson, 1999; Goodhand, 2002; Lischer, 2003; Maren, 2009; Polman, 2010; Barnett, 2013; Zürcher, 2019; O’Donnell, 2023) and cross-national inquiries also suggest that aid prolongs conflict in part due to diversion (Nunn and Qian, 2014; Narang, 2015; Wood and Sullivan, 2015; Wood and Molfino, 2016; Findley, 2018; Findley et al., 2023). (See Table A10, Column 8).
@@ 3 Costs of Delivery
The last three columns of Table A10 report whether the programs (a) reported on and delivered aid cost-efficiently or cost-effectively, (b) decentralized their aid distribution, and (c) could transparently track aid receipt and usage.
@@
Of the 23 reviewed programs, 12 report cost-efficiency or cost-effectiveness metrics. (Table A10 reports the derivable total cost transfer ratios (TCTRs) for nine of the 12; the remaining three, which we tag with a check mark (✓), either reported on bundled programs for which we could not disaggregate the costs or did not include the precise estimates in their reports.) The cost efficiency of our digital aid program is the lowest of any of these programs when comparably counting cost categories. Although all of the reviewed studies focus on variable operational costs, they vary in the specific cost categories they count as relevant. In particular, very few studies include targeting/recruitment costs, and are thus comparable to our TCTR of \$1.014 (1.4 cents required to deliver 1 dollar of aid), though we think the inclusion of targeting/recruitment costs is more appropriate (and conservative) given that these costs are likely to be present in most crisis contexts.
@@ 1 Research design
Afghanistan is in its third year of a complex humanitarian crisis combining economic contraction, political repression, and ongoing concerns over stability and security. Its humanitarian appeal of \$4.6 billion in 2023 represented the world’s largest single-country appeal to date (United Nations, 2023c). It is also the WFP’s largest recipient by value of cash-based transfers and commodity vouchers (World Food Programme, 2023a) and though multiple financial service providers exist in the country, digital aid programs are still nascent (United Nations Capital Development Fund, 2023). Afghanistan’s GDP contracted by 30-35% since the Taliban took over in 2021, and ever since the country has been under the threat of “winters of famine" (World Bank, 2023a; World Food Programme, 2023d).
@@
Table A5 reports summary statistics for the experimental sample. Most participants are widows (66%), have no education (63%) and are the main financial decision-maker in the household (66%), which on average has 6.31 members, indicating that the women in the study are highly vulnerable. Using the 2015 Demographic and Health Survey (DHS), we compare our sample to a representative sample of similarly-aged women in urban areas of Kabul, Balkh and Mazar. The DHS sample has higher educational attainment: 56.8% have no schooling and 13.9% have at most primary education.
@@
A team of 12 female enumerators were tasked with contacting each participant once every month. Participants were randomly assigned a date to be contacted. Enumerators contacted participants on said date, and if the survey was not completed (because the participant did not pick up or was busy), they would attempt again on a different date and time. Participants received a 350 AFA payment as compensation for their time for completing their survey. Overall, for the two survey rounds that correspond to the experimental sample, completion rates were around 99%. In the first survey round, 29 interviews could not be completed (17 treatment, 12 control). In the second round, 26 interviews could not be completed (10 treatment, 16 control). Attrition is not differential by treatment status, as seen in Table A12, but there is a slightly higher non-response rate among control participants in the second round of surveys (the difference is 6 surveys).
@@ 2 Analysis
Given that participants were allowed to skip certain questions if they wanted, this means that the sample in Table 1 changes in each regression. We also provide results restricting the sample to only those individuals who answered all questions necessary to construct our primary and secondary outcomes, ensuring a constant sample across regressions. These results are presented in Table A4, which follows the same structure as Table 1. Results using this alternative approach are consistent with the baseline results in Table 1.
@@
where $\mathbb{1}[\textit{Heterogeneity}]_{in}$ is a dummy for whether woman $i$ in Nahia $n$ belongs to a certain subgroup. We check for heterogeneity along the following dimensions: basic needs (total meals composed of only bread and tea), city of residence (Kabul vs. Mazar and Herat), whether the woman is able to leave the house at baseline, whether the woman is married, whether the woman is Pashtun, whether the woman has some education, whether the woman is above the median age, whether the woman is the household’s financial decision-maker, and whether the household is above the median household size. For brevity, we only present results for the KLK indices of the three main outcome families in Figure A3.
@@
Results of this exercise are presented in Figure A4. Even with a fee as small as 100 AFN (2.5% of the total payment), almost 80% prefer receiving the aid digitally rather than in cash. A fee of 300 AFN that more accurately reflects the true costs of delivering the aid in cash results in 94% of respondents preferring digital aid. The treatment group has a slightly higher preference for digital aid than the control group, potentially because they have had more experience receiving the aid digitally, although responses are very similar across the two groups.
@@ 3 Identification Threats
The results of the experimenter demand effect analysis are shown in A3. Column 2 shows the baseline estimates, $\hat{\gamma}_{1}$, column 3 shows the estimates for the overall experimenter demand effects, $\hat{\psi}_{1}$, column 4 shows the experimenter demand effects only for the control group, $\hat{\eta}_{1}$, and column 5 shows the (overall) experimenter demand effects for the treatment group, $\hat{\eta}_{1}+\hat{\eta}_{2}$.
@@ 4 Deviations from PAP
As mentioned before, our PAP is registered here. All the analysis was conducted after the submission of this PAP.
@@
We adhered to the PAP as closely as possible, although there were a few instances in which we deviated. We do not report Treatment on the Treated (ToT) estimates because there were no issues with non-compliance and survey response rates were extremely high. Given that attrition was so low and not different across treatment groups, we also do not show Lee Bounds results. We do not report results on participants’ satisfaction with the digital payments platform, HesabPay, but there were only 2 instances of participants stating that they were “somewhat dissatisfied” with HesabPay (out of 2393 responses in the treatment group), with 2358 reporting being “very satisfied”. In addition to testing for experimenter demand effects for the whole primed sample (pre-specified), we also show results disaggregated by treatment group (not pre-specified). We also don’t report results on the “Response timing” section of the PAP, as there were no significant results (results are available upon request). Initially, we had described a broader set of questions for the experts’ survey, but we decided to cut the number of questions from six to four in order to reduce the burden on respondents. The four questions we did ask match the topics in the PAP. We created a table summarizing all the main (pooled) results, in which the sample was restricted to the participants who answered all the relevant questions across the two months of surveying, to allay concerns of comparability (see Table A4). As a robustness check, we also include an alternative way of creating summary indices following (Anderson, 2008), another popular approach for constructing indices. This was also not pre-specified.
@@
We deviate slightly from the PAP in the diversion results. Initially, we had pre-specified presenting results for each of the four measures of diversion, as well as an index created by combining the four measures following (Kling, Liebman, and Katz, 2007), the KLK index. Upon further consideration, and after computing the results, we decided to additionally show results using a dummy for whether the respondent answered yes to any of the four diversion tables (column 6 in Table 3), which was not pre-specified. This is because, due to the lack of positive responses to the informal taxation questions and the way in which the index is created (standardising first each component variable, and then the final index) leads to extremely large values for those who answered yes to multiple questions. For reference, out of the 59 instances of participants answering yes to any of the four questions in any of the two survey rounds, 49 said yes to only one question, 9 said yes to two questions, and only one said yes to three questions (e.g. index value for this woman is 34, a clear outlier), with no one saying yes to all four. For completeness and robustness, Table A6 shows the results in which the two measures of diversion of others are combined into a single KLK index (column 1), the two measures of diversion of the participants themselves are combined into a single KLK index (column 2), and an index of the four measures together following (Anderson, 2008) (column 3). Moreover, we also present results dropping the one outlier that answered yes to three of the taxation questions in one round and received a KLK Index value above 34 standard deviations (column 4). None of these alternative indices was pre-specified. As can be observed, while the results using the pre-specified KLK index in Table 3 are only significant at the 10% level, the other four alternative indices are all statistically insignificant at conventional levels. This suggests that the one outlier is what leads to the significant results in the main results. We also do not report the $p$-values adjusted for multiple hypothesis testing for the individual diversion measures, but they are all above 0.25.

### v4

@@ 1 Cash-based transfers in humanitarian contexts
In all, we identified 23 relevant cash-based transfer programs summarized in A10. Peer-reviewed academic studies account for 16 of the 23, five are academic working papers, and two are evaluation reports. Column 1 reports the country in which the cash-based intervention took place and the countries are ordered by the current level of fragility of the country in which they occur, moving from the least fragile (top) to most fragile (bottom) (OECD, 2022). Ecuador and Philippines appear at the top because neither country is listed in the OECD’s list of fragile countries. The next two, Lebanon and Sri Lanka, are not on the core OECD list either, but are flagged as showing key early-warning signs. Sometimes the target beneficiaries (Column 2) were refugees from another country, but located in a host country, and sometimes the intended beneficiaries were local to a country. Regardless, target beneficiaries were high need or otherwise vulnerable. Column 3 reports the type of cash-based assistance, specifically identifying cash (C) and vouchers (V), and the most relevant alternative, food assistance (F). Some of these programs include other/combined treatments, which we note in the extended discussions. Column 4 reports the delivery mechanism, whether some form of digital, physical, or a combination.
@@ 2 Work related to outcomes: Usage, food security, well-being, diversion
To contextualize our experimental findings on usage, food security and mental well-being, A10 reports whether the 23 reviewed studies empirically examined usage at any level, denoting examination with a check mark (✓), the same (or similar) outcome categories that we did, including whether they found positive, null, or negative effects. Few studies carried out systematic tests for diversion, so we only include a check mark (✓) if the study made any observation about levels of diversion in their context. To supplement the summary table, LABEL:LiteratureFindings1 provides details about specific study context and findings.
@@
Because tech-illiterate populations are largely poor, rural, elderly, and female (Peachey, 2020), it is important to assess whether digital transfers will reach those most vulnerable and be used effectively. If they cannot, then digital cash-based approaches may hurt intended beneficiaries relative to what may have occurred with physical cash or in-kind food distribution. In India’s PMGKY program intended to provide COVID relief, for example, the government transferred benefits to the accounts of over 200 million beneficiaries, but being female, illiterate, and living in a household without a smartphone resulted in extremely low use of digital payments, which ranged from 1% to 3.9% (Gentilini, 2022). Of the 10 studies in our review with a digital component, five relied at least in part on mobile platforms, meaning that the beneficiary needed to use the phone for receipt/withdrawal of funds. The remaining five relied primarily on ATM cards, which are less demanding technologically, and also confer fewer advantages such as decentralization and transparency. Of these five, only one reported on usage, noting that most of the time the funds were withdrawn immediately and, where not, they were used up eventually (Lyall, Zhou, and Imai, 2020). The remaining four did not report usage explicitly, but descriptions were suggestive of substantial usage, though we acknowledge that this assessment may be overly generous. Our own results in one of the most challenging environments, and with an arguably more vulnerable demographic than most other studies, are encouraging, though we emphasize the importance of user-centered design in our registration and training procedures. Near 100% usage by a largely illiterate and vulnerable population in a context of extreme oppression and insecurity suggests genuine potential for technology to “flatten access” to otherwise marginalized populations (Grossman, Humphreys, and Sacramone-Lutz, 2014). (See A10, Column 5).
@@
Most of the studies (21 of 23) considered a food security outcome in some form, although there is considerable variation in the indicators each examines. Of the 21, 16 find that cash-based transfers improved food security outcomes, five found no relationship, and none of the studies found any negative effects. Of the 21, 20 examined at least one indicator of food quantity (e.g., food expenditures or calorie consumption). Of the 21, 18 examined at least one indicator of food quality (e.g., dietary diversity). And 16 of 21 considered at least one indicator of negative coping strategies (e.g., meal skipping). These patterns add support to a growing body of evidence documenting that humanitarian cash-based programs have positive impacts on a variety of food security outcomes. Existing studies most frequently considered basic needs outcomes. In a recent systematic review, eight of 20 studies considered food security (primarily food consumption) and four of 20 studies examined dietary diversity, most of which demonstrated positive impacts (Jeong and Trako, 2022). Moreover, five studies examined reductions in negative coping strategies (e.g., meal skipping) and three found positive impacts. (See A10, Column 6).
@@
In contrast to food security outcomes, A10 shows that limited attention has been given to the possible mental benefits (or costs) of cash-based programs. Only seven of the 23 studies considered some mental well-being outcomes and six found positive effects. An evaluation of a voucher program in DRC found a positive impact (+0.32sd) on an overall mental well-being index at the end of six weeks, which attenuated to some extent, but remained remarkably durable at the end of one year (+0.18sd and statistically significant) (Quattrochi et al., 2022). They also measured life satisfaction, where they found a positive impact at six weeks (18% higher in treatment; control: 3.29 on 10-pt scale), but the result disappeared by the one-year mark. Perhaps the strongest impact of cash on mental well-being, with documented short-, medium-, and long-term effects of cash relative to control, occurred in Niger (Bossuroy et al., 2022). Importantly, like the other treatment arms, cash was accompanied by coaching, savings, training, and market facilitation, making the overall bundle difficult to compare to other cash-only programs. The dearth of studies examining mental well-being mirrors that of prominent systematic reviews (Jeong and Trako, 2022). They only identified two (of 20) studies reporting subjective well-being results: the DRC voucher study that we just discussed (Quattrochi et al., 2022) and a study of Syrian refugees in Jordan (that did not meet our inclusion criteria) that found isolated improvement on a single measure of self-esteem, but notably no effect on a measure of satisfaction, the closest to our life satisfaction indicator. Similarly, in another systematic review, there were no rigorously evaluated studies that addressed mental well-being though some mixed-methods evaluations have found isolated positive impacts (Attah et al., 2016; Doocy and Tappis, 2016; Hagen-Zanker et al., 2017). (See A10, Column 7).
@@
The extent of diversion in these studies is essentially unknown as four of the 23 studies addressed diversion, but only three attempted to measure its extent. Studies measured the amount of the cash and vouchers actually received and found little leakage (Aker, 2017), how much of the grant recipients had to give to other household and community members and participants who reported less than 1% of the grant (Blattman et al., 2016), and the occurrence of robbery where participants reported no occurrence (Lehmann and Masterson, 2020). Another study mentioned that the implementer feared diversion, and therefore hired security to guard cash during transport, which comprised the largest share of their costs, but did not report measuring diversion (Aker et al., 2016). Many policy reports have downplayed concerns about diversion, suggesting that diversion risks are likely overstated and that cash is no less prone to diversion than other types of aid (Sossouvi, 2013; Gordon, 2015; Overseas Development Institute, 2015; Ecker, Maystadt, and Guo, 2019). As far as we are aware, no systematic tests of diversion in a cash-based context have been undertaken. In our study, we experimentally tested for diversion, both using direct and indirect questions, and found no evidence of its presence. We even interviewed key stakeholders, such as merchants, to understand alternative points where diversion could have occurred, and found little indication. This finding is especially informative given how extensively the scholarly and policy literatures have identified diversion as a core concern for in-kind assistance. Indeed, anecdotal evidence of diversion has been extensively documented across the globe (Perlez, 1992; Bryer and Cairns, 1997; De Waal, 1997; Anderson, 1999; Goodhand, 2002; Lischer, 2003; Maren, 2009; Polman, 2010; Barnett, 2013; Zürcher, 2019; O’Donnell, 2023) and cross-national inquiries also suggest that aid prolongs conflict in part due to diversion (Nunn and Qian, 2014; Narang, 2015; Wood and Sullivan, 2015; Wood and Molfino, 2016; Findley, 2018; Findley et al., 2023). (See A10, Column 8).
@@ 3 Costs of Delivery
The last three columns of A10 report whether the programs (a) reported on and delivered aid cost-efficiently or cost-effectively, (b) decentralized their aid distribution, and (c) could transparently track aid receipt and usage.
@@
Of the 23 reviewed programs, 12 report cost-efficiency or cost-effectiveness metrics. (A10 reports the derivable total cost transfer ratios (TCTRs) for nine of the 12; the remaining three, which we tag with a check mark (✓), either reported on bundled programs for which we could not disaggregate the costs or did not include the precise estimates in their reports.) The cost efficiency of our digital aid program is the lowest of any of these programs when comparably counting cost categories. Although all of the reviewed studies focus on variable operational costs, they vary in the specific cost categories they count as relevant. In particular, very few studies include targeting/recruitment costs, and are thus comparable to our TCTR of \$1.014 (1.4 cents required to deliver 1 dollar of aid), though we think the inclusion of targeting/recruitment costs is more appropriate (and conservative) given that these costs are likely to be present in most crisis contexts.
@@ 1 Research design
Afghanistan is in its third year of a complex humanitarian crisis combining economic contraction, political repression, and ongoing concerns over stability and security. Its humanitarian appeal of \$4.6 billion in 2023 represented the world’s largest single-country appeal to date (United Nations, 2023c). It is also the WFP’s largest recipient by value of cash-based transfers and commodity vouchers (World Food Programme, 2023a) and though multiple financial service providers exist in the country, digital aid programs are still nascent (United Nations Capital Development Fund, 2023). Afghanistan’s GDP contracted by 30-35% since the Taliban took over in 2021, and ever since the country has been under the threat of “winters of famine” (World Bank, 2023a; World Food Programme, 2023d).
@@
A6 reports summary statistics for the experimental sample. Most participants are widows (66%), have no education (63%) and are the main financial decision-maker in the household (66%), which on average has 6.31 members, indicating that the women in the study are highly vulnerable. Using the 2015 Demographic and Health Survey (DHS), we compare our sample to a representative sample of similarly-aged women in urban areas of Kabul, Balkh and Mazar. The DHS sample has higher educational attainment: 56.8% have no schooling and 13.9% have at most primary education.
@@
A team of 12 female enumerators were tasked with contacting each participant once every month. Participants were randomly assigned a date to be contacted. Enumerators contacted participants on said date, and if the survey was not completed (because the participant did not pick up or was busy), they would attempt again on a different date and time. Participants received a 350 AFA payment as compensation for their time for completing their survey. Overall, for the two survey rounds that correspond to the experimental sample, completion rates were around 99%. In the first survey round, 29 interviews could not be completed (17 treatment, 12 control). In the second round, 26 interviews could not be completed (10 treatment, 16 control). Attrition is not differential by treatment status, as seen in A12, but there is a slightly higher non-response rate among control participants in the second round of surveys (the difference is 6 surveys).
@@ 2 Analysis
Given that participants were allowed to skip certain questions if they wanted, this means that the sample in Table 1 changes in each regression. We also provide results restricting the sample to only those individuals who answered all questions necessary to construct our primary and secondary outcomes, ensuring a constant sample across regressions. These results are presented in A5, which follows the same structure as Table 1. Results using this alternative approach are consistent with the baseline results in Table 1.
@@
where $\mathbb{1}[\textit{Heterogeneity}]_{in}$ is a dummy for whether woman $i$ in Nahia $n$ belongs to a certain subgroup. We check for heterogeneity along the following dimensions: basic needs (total meals composed of only bread and tea), city of residence (Kabul vs. Mazar and Herat), whether the woman is able to leave the house at baseline, whether the woman is married, whether the woman is Pashtun, whether the woman has some education, whether the woman is above the median age, whether the woman is the household’s financial decision-maker, and whether the household is above the median household size. For brevity, we only present results for the KLK indices of the three main outcome families in A2.
@@
Results of this exercise are presented in A3. Even with a fee as small as 100 AFN (2.5% of the total payment), almost 80% prefer receiving the aid digitally rather than in cash. A fee of 300 AFN that more accurately reflects the true costs of delivering the aid in cash results in 94% of respondents preferring digital aid. The treatment group has a slightly higher preference for digital aid than the control group, potentially because they have had more experience receiving the aid digitally, although responses are very similar across the two groups.
@@ 3 Identification Threats
The results of the experimenter demand effect analysis are shown in A4. Column 2 shows the baseline estimates, $\hat{\gamma}_{1}$, column 3 shows the estimates for the overall experimenter demand effects, $\hat{\psi}_{1}$, column 4 shows the experimenter demand effects only for the control group, $\hat{\eta}_{1}$, and column 5 shows the (overall) experimenter demand effects for the treatment group, $\hat{\eta}_{1}+\hat{\eta}_{2}$.
@@ 4 Deviations from PAP
As mentioned before, our PAP is registered at
https://www.socialscienceregistry.org/versions/160809/docs/version/document. All the analysis was conducted after the submission of this PAP.
@@
We adhered to the PAP as closely as possible, although there were a few instances in which we deviated. We do not report Treatment on the Treated (ToT) estimates because there were no issues with non-compliance and survey response rates were extremely high. Given that attrition was so low and not different across treatment groups, we also do not show Lee Bounds results. We do not report results on participants’ satisfaction with the digital payments platform, HesabPay, but there were only 2 instances of participants stating that they were “somewhat dissatisfied” with HesabPay (out of 2393 responses in the treatment group), with 2358 reporting being “very satisfied”. In addition to testing for experimenter demand effects for the whole primed sample (pre-specified), we also show results disaggregated by treatment group (not pre-specified). We also don’t report results on the “Response timing” section of the PAP, as there were no significant results (results are available upon request). Initially, we had described a broader set of questions for the experts’ survey, but we decided to cut the number of questions from six to four in order to reduce the burden on respondents. The four questions we did ask match the topics in the PAP. We created a table summarizing all the main (pooled) results, in which the sample was restricted to the participants who answered all the relevant questions across the two months of surveying, to allay concerns of comparability (see A5). As a robustness check, we also include an alternative way of creating summary indices following (Anderson, 2008), another popular approach for constructing indices. This was also not pre-specified.
@@
We deviate slightly from the PAP in the diversion results. Initially, we had pre-specified presenting results for each of the four measures of diversion, as well as an index created by combining the four measures following (Kling, Liebman, and Katz, 2007), the KLK index. Upon further consideration, and after computing the results, we decided to additionally show results using a dummy for whether the respondent answered yes to any of the four diversion tables (column 6 in Table 3), which was not pre-specified. This is because, due to the lack of positive responses to the informal taxation questions and the way in which the index is created (standardising first each component variable, and then the final index) leads to extremely large values for those who answered yes to multiple questions. For reference, out of the 59 instances of participants answering yes to any of the four questions in any of the two survey rounds, 49 said yes to only one question, 9 said yes to two questions, and only one said yes to three questions (e.g. index value for this woman is 34, a clear outlier), with no one saying yes to all four. For completeness and robustness, A7 shows the results in which the two measures of diversion of others are combined into a single KLK index (column 1), the two measures of diversion of the participants themselves are combined into a single KLK index (column 2), and an index of the four measures together following (Anderson, 2008) (column 3). Moreover, we also present results dropping the one outlier that answered yes to three of the taxation questions in one round and received a KLK Index value above 34 standard deviations (column 4). None of these alternative indices was pre-specified. As can be observed, while the results using the pre-specified KLK index in Table 3 are only significant at the 10% level, the other four alternative indices are all statistically insignificant at conventional levels. This suggests that the one outlier is what leads to the significant results in the main results. We also do not report the $p$-values adjusted for multiple hypothesis testing for the individual diversion measures, but they are all above 0.25.